The power required to push a car is 1696W(watts).
After the car is pushed fully 3696J work is done.
Find the time it took to push the car.
Answer #1

given

P = 1696 W

W = 3696 J

t = ?

P = W / t

t = W / P

= 3696 / 1696

t = 2.179 sec or 2.18 sec

so the time it took to push the car is t = 2.179 sec or 2.18 sec